OpenStudy (anonymous):

Diana made a recipe that yields 6 and 1 over 2 cups. If each serving is 1 over 4 cup, which expression will help Diana determine the number of servings her recipe will yield?

OpenStudy (anonymous):

Diana made a recipe that yields 6 and 1 over 2 cups. If each serving is 1 over 4 cup, which expression will help Diana determine the number of servings her recipe will yield? 6 and 1 over 2 ⋅ 1 over 4 6 and 1 over 2 ÷ 1 over 4 6 and 1 over 2 + 1 over 4 6 and 1 over 2 – 1 over 4

OpenStudy (anonymous):

can anyone help??

OpenStudy (anonymous):

6 and 1 over 2 is mathematically written as \[6\frac{ 1 }{ 2 }\] and to write this as a simple fraction, you multiply the denominator with the whole number and add the numerator. the result will be the new numerator and the denominator remains the same as before. Being \[\frac{ 2*6+1 }{ 2 }=\frac{ 13 }{ 2 }\]. If one serving takes \[\frac{ 1 }{ 4 } \] of the recipe, then you need to find x. becoming \[\frac{ 13 }{ 2 }=\frac{ 1 }{ 4 }x\] Find x.

OpenStudy (anonymous):

how do I find x?

OpenStudy (anonymous):

@Hoslos

OpenStudy (anonymous):

You have to transport the fraction on the right hand side to the left one, but since it was being multiplied by x, it will be divided by 13/2 in the left hand side. Plus, remember that in division of fraction, you multiply the first one with the inverse of the second.

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

B. What I told you is B. It is division with 1/4, not multiplication.